Circular No. 6747

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 A. Fruchter, M. Livio, D. Macchetto, L. Petro, and K. Sahu, Space

Telescope Science Institute; E. Pian and F. Frontera, ITESRE, CNR, Bologna; S. Thorsett, Princeton University; and M. Tavani, Columbia University and IFCTR, Milan, report: “The optical counterpart of GRB 970228 (cf. IAUC 6584, 6588; van Paradijs et al. 1997, Nature 386, 686) was imaged via eight exposures of 575 s with the Hubble Space Telescope (HST) STIS CCD (+ clear filter) between Sept. 4.65 and 4.76 UT. Both the nebula and point-source optical transient found by Sahu et al. (loc. cit. 387, 476) in earlier HST WFC2 observations are detected in the STIS images. The nebula is roughly circular with an apparent diameter of ∼ 0′′.8, and the optical transient lies within the nebula but near its southern edge. We determine V = 28.0 ± 0.25 for the point source and V = 25.7 ± 0.15 for the nebula. Reanalysis of the earlier HST WFPC2 F606W observations of this source taken on Mar. 26 and Apr. 7, using an aperture determined from the STIS image, gives V = 25.6 ± 0.25 for the nebula. The diameter of the nebula is comparable to galaxies of similar magnitude found in the Hubble Deep Field, especially if one takes into account a possible visual extinction in the direction of GRB 970228 of at least one magnitude (D. Lamb, Huntsville Symposium). The position of the point source in the STIS images shows a proper motion with respect to earlier HST images of 25 ± 15 mas (see also IAUC 6629); we interpret this as a 2σ upper limit to any proper motion of ∼ 100 mas per year. The temporal decay of the optical transient appears to be well-fitted by a power-law with index of ∼ −1.1, with the exception of data points taken between Mar. 6 and 13, all of which fall below the power-law (see Galama et al., loc. cit. 387, 479). However, the long-term agreement between the GRB lightcurve and the power law supports the relativistic fireball model, and thus implies that GRB 970228 is an extragalactic phenomenon.”

COMET 103P/HARTLEY 2 Total visual magnitude estimates: Aug. 30.02 UT, 13.9 (J. Carvajal,

Avila, Spain, 0.32-m reflector); Sept. 5.79, 14.9 (G. Bakos, Ráktanya, Hungary, 0.30-m reflector); 7.87, 13.2 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 8.84, 13.6 (W. Hasubick, Buchloe, Germany, 0.44m reflector); 21.82, 12.6 (Bouma).

1997 September 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green